OpenStudy (anonymous):

find in terms of k the sum of all the odd integers lying btwn 2k and 4k

OpenStudy (anonymous):

We have that the number of terms is (4k-2k)/2 =k, the first term is 2k +1 and last term 4k-1 so using the formula S = n/2 (a_1 +a_n) S = k/2 (2k +1 + 4k -1) = k/2(6k) = 3k^2.

OpenStudy (anonymous):

how did you get the number of terms part?

OpenStudy (anonymous):

Well, take any 2 actual numbers of form 2k, say 40 and 36, the number of numbers after 36 is 4 but we only want half of them so 40-36/2

OpenStudy (anonymous):

i dont understand.......=(

OpenStudy (anonymous):

There's two kinds of integers, odd and even and they alternate.

OpenStudy (anonymous):

Is it something else u don't understand?